Uncartainty in AI - Autumn 1996

SEMINAR TALK SCHEDULE

Date  Who  What
November 10  Cohavit Taboch and Samir Genaim Statistical NLP
November 17  Carmel Dumshlak  Hierarchical planning in stochastic domains
November 24  Dimitri Yukelson  Neural Networks, introduction
December 1  Genadi Polishuk  Case-Based Reasoning
December 8  No talks (Hannukah)
December 15  Mark Kharitonov  Neural Networks
December 22  Nickolai Berezanski  Decesion-theoretic Search Control
December 29  Merav Cohen  Probabilistic Rule-Based Systems
January 5, 1997 Yevgeny Schechtman  Probabilistic Logic
January 12  Shalom Oren?  ???